# Accessing the topological properties of human brain functional sub-circuits in Echo State Networks

### v1

@@ 1 Introduction
One of the prominent directions in building high-performance or general intelligence artificial systems is implementing biologically plausible models, particularly using human brain connectomes as the underlying graph of neural networks Damicelli et al. (2022); Suarez et al. (2020); McDaniel et al. (2022); Suarez et al. (2023). Most state-of-the-art biologically inspired artificial connectomes are realized as reservoir networks with structural human brain data. However, structural connectomes (SC) derived from diffusion Magnetic Resonance Imaging (MRI) are generally rigid in both topology and scale compared to functional connectomes (FC) from functional MRI data. Thus, in this work, we explore the prospect of applying functional connectomes to reservoir Echo-state Networks (ESNs) under various conditions to comprehensively evaluate model performance, as well as build a consistent framework for implementations of connectomic data in artificial networks. Motivated by the fact that the topology of a neural network has a tangible impact on ESN performance, and that complex network topology sometimes improves the performance of the model, we systematically explore more complex and biologically inspired topology in reservoir neural networks. A critical step in ESN embedding requires pre-determined read-in and read-out layers with inputs sampled from the induced subgraphs of the embedded reservoir. Typically, a priori set of functional networks (FNs), such as Yeo’s FNs Yeo et al. (2011), is used as pre-determined subgraphs for read-in and read-out input selections. Since a priori set of FNs are derived from functional MRIs, yet, these pre-determined sub-circuits are mapped onto SC, it is unknown whether FN-to-SC embedding is well justified from i) neuro-physiological perspective and ii) ESN performance. In this paper, our contributions are as follows: i) We proposed a pipeline to implement and evaluate echo-state reservoir networks with various embedded topology on various processing and memorization tasks. To this end, we tested various topological embedding and evaluated their respective performance on diverse tasks, showing different performance optimums are highly dependent on the embedded neuro-physiological architecture of a priori functional networks and the corresponding task configurations; ii) we show that, in general, complex topology perform better than simple bipartite models or null models with feed-forward properties commonly seen in MLP for certain tasks and reservoir criticality conditions. As such, we also found that, contrary to earlier literature, the performance of the reservoir model is not strictly defined by the reservoir’s echo-state property defined by the spectral radius of the connectivity matrix; iii) we found that there exists a differentiated performance of functional networks induced from the human structural connectome with respect to different cognitive tasks and their corresponding computation-memorization balances. We analyzed the topological properties of a priori functional networks and highlighted their graph-theoretical properties that play significant roles in determining the final performance of the echo-state model.
@@ 2.1 Topologically-Embedded artificial neural network and Echo-State-Networks (ESN)
Echo-state networks are a variant of the classical reservoir computing (RC) paradigms, where the reservoir is a recurrent layer of the classical recurrent neural network Verstraeten et al. (2007). RCs are typically proposed in place of regular RNN variants as a lightweight model, where the hidden states are high dimensions while only the readout layer weights of the model must be updated during training. The frozen nature of the reservoir layer allows any topology and unit dynamics to be embedded. As a result, RCs have been implemented as predictors of chaotic dynamics in physics simulators, or used as a time-series prediction model in general Chattopadhyay et al. (2020); Shahi et al. (2022); Bompas et al. (2020); Huhn & Magri (2022); Platt et al. (2022). ESNs model time as discrete steps, making them suitable for discrete-time sampling data seen in real-world time-series datasets Cucchi et al. (2022). In the current literature, the relationship between the topological structure of an NN and its performance is still unclear. While complex networks may yield higher predictive performance or parameter efficiency Kaviani & Sohn (2021), it is difficult to generalize randomly wired models outside the investigated context. Under the constraint of MLP-random interpolation, several random network models rewired similarly to the small-world regimes perform well on several real-life problems Erkaymaz et al. (2017); Erkaymaz & Ozer (2016). Different topology perform well in limited tasks on small network sizes enabled by the learning matrix sequential algorithm for training directed acyclic graphs (DAGs). Random graph topology are also shown to perform well as image classifiers on the ImageNet dataset, especially the Watts-Strogatz (WS) model of small-world networks Xie et al. (2019). Boccato et al. (2024) echoes the statement on the performance of small-world models in the case of synthetic data, albeit the difference between WS and other complex graph families is less pronounced, possibly owing to the sufficient degree of topological complexity for function representation.
@@ 2.2 Brain Connectomics and human brain functional sub-circuits
Human Brain connectomics studies how the human brain is structurally connected Xu et al. ; Xu et al. (2022; 2024); Garai et al. (2024b) and how a heterogeneous repertoire of resting-state or task-related functional circuits emerge on top of it Abbas et al. (2020; 2023); Amico et al. (2019; 2021); Chiêm et al. (2022); Duong-Tran et al. (2024a; c; b; d); Nguyen et al. (2024); Garai et al. (2023; 2024b; 2024a). There are two types of mesoscopic structures in large-scale brain networks: localized and non-localized. In this paper, we focus on investigating localized mesoscopic structures which are sub-systems learned from local/quasi-local network properties such as brain regions of interest (ROIs) or functional edges, or correlations among neighboring nodes Duong-Tran et al. (2024a). In brain connectomics, these sub-structures are induced from a wide array of techniques, including but not limited to clustering Yeo et al. (2011); Power et al. (2011) or low-dimensional approximation of high-dimensional dynamics Shine et al. (2016); Shine & Poldrack (2017); Shine et al. (2019). The most commonly known localized mesoscopic structures in brain networks are often referred to as functional sub-circuits or functional networks (FNs) Yeo et al. (2011); Sporns & Betzel (2016); Duong-Tran et al. (2019; 2024a). From here on, we will use the two terms functional sub-circuits and FNs interchangeably.
@@ 2.3 Human-Brain-Connectomics ESNs
In the context of modeling the human brain connectome, ESNs are implemented as prototype models of the structural brain topology. Suarez et al. (2020) shows that human brain connectivity captured through diffusion imaging performs better than random null network models in the critical regime, while also demonstrating computationally relevant properties of resting-state functional brain network parcellation. Given a minimum level of randomness and connection diversity from the original structural connectome, biologically inspired reservoir networks’ performance matches that of any classical random network model Damicelli et al. (2022). The authors of d’Andrea et al. (2022) conclude that the modular structure of reservoirs significantly impacts prediction error among multiple other features. While the connections are modeled in a recurrent network rather than strictly an RC, Achterberg et al. (2023) shows that training an RNN constrained by reservoir features such as spatial wiring cost and neuron communicability leads to the emergence of structural and functional features commonly found in primate cerebral cortices.
@@ 3.1 Preliminary: Echo-state Implementation
To examine different topologies as computational graphs, we utilized the implementation of the classical ESN Jaeger (2001b) from the conn2res package Suarez et al. (2023) which has the following form:
@@
where $\textbf{X}\in\mathds{R}^{N_{p}\times T}$ is the matrix containing temporal reservoir states $x[t]$ of a certain $N_{p}$ nodes ($N_{p}<N_{r}$) of the reservoir computed from the previous states and the input $u[t]$ for $t=1,...,T$, $\textbf{Y}\in\mathds{R}^{N_{o}\times T}$ is the ground-truth matrix containing $y[t]$, $\textbf{I}\in\mathds{R}^{N_{p}\times N_{p}}$ is the identity matrix, and $\lambda$ is the regularization parameter. Additionally, we keep track of the spectral radius $\alpha$ of the reservoir’s adjacency matrix to examine the reservoir’s adherence to the echo-state property (ESP), a condition quantifying the reservoir’s unique stable input-driven dynamics Jaeger (2001b).
@@ 3.2 Pipeline and structural control model
The reservoir network is generally partitioned using intrinsic functional networks, a connectivity-based partition of brain network into functionally similar groups of areas Wig (2017). The reservoir input weight matrix $\textbf{W}^{I}\in\mathds{R}^{n\times m}$ route the input signals to only certain $m$ nodes in the reservoir (Figure 1, 2). The input nodes are chosen as an entire functional network to examine message propagation on the reservoir computational graph; in this paper, we chose subcortical regions as input nodes, owning to the plausibility of these regions serving as relay stations for incoming sensory signals. Each of the other seven cortical regions is chosen as the read-out node-set, where the readout weight matrix $\textbf{W}^{O}$ connects the $N_{p}$ nodes of the functional network to the readout module.
Figure 2: Human brain functional networks (sub-circuits) as parcellated by Yeo and colleagues Yeo et al. (2011). Figure is referenced from Duong-Tran et al. (2021).
@@
We embed the structural human brain connectome into the reservoir as the default topology map. Consensus adjacency matrices are constructed from the original connectomes with bootstrap resampling to provide a reliable estimate of the model performance. The structural connectome is grouped into seven intrinsic functional networks and one subcortical region according to the Yeo functional mapping and parcellation Thomas Yeo et al. (2011). These structures are known as sub-circuits or FNs, previously defined in Section 2.2.
@@ 3.3 Variants: Complex graph null models
We preserve the original degree-sequence of the structural connectome and rewired each edge in the connectome exactly once using the Maslov-Sneppen algorithm Maslov & Sneppen (2002). From the rewiring procedure, we obtained a reference null connectome with the original constraint still moderately enforced.
@@
An entirely new graph is constructed from the degree sequence of the original structural connectome, as opposed to only rewiring each edge once. The configuration model generates a random undirected pseudograph by randomly assigning edges to match a given degree sequence Newman (2003). Self-loops are then removed from the existing graph to obtain the symmetric null connectome adjacency matrix for the reservoir.
@@ 4.1 Data
Adapted from structural consensus connectomes from Suarez et al. (2020). The connectome is divided into 463 or 1015 approximately equally sized nodes. A group-consensus approach is adopted to mitigate reconstruction inconsistencies Betzel et al. (2019) and network measure sensitivities between different maps. Functional connectomes are obtained from a publicly available dataset (Derived Products from HCP-YA fMRI, Tipnis et al. (2021)). Structural networks are taken from Suarez et al. (2020) where 66 participants were scanned using diffusion MRI under no task demands, divided into 463 or 1015 approximately equally sized nodes. The dataset contains fMRI time-series parcellated using the Schaefer parcellation at resolutions from 100 to 900 parcels; we use the parcellation resolution at 500 nodes via the Yeo 7-region functional networks schema without Global Signal Regression (GSR) under resting state.
@@
We used synthetic behavioral data to simulate computation from multiple signal sources and assess the memory capacity of the reservoir under various means. The PerceptualDecisionMaking (PDM) Britten et al. (1992) task is a two-alternative forced choice task requiring the model to integrate and compare the average of two stimuli, with the PerceptualDecisionMakingDelayResponse (PDMDR) Inagaki et al. (2019) variant artificially inducing a memory requirement by prompting the model to answer after a random delay. The context-dependent decision-making task ContextDecisionMaking (CDM) Mante et al. (2013) requires the model to make a perceptual decision based on only one of the two stimulus inputs from two different modalities, where a rule signal indicates the relevant modality. The assessment for the memory capacity of the model is the MemoryCapacity (MemCap) Jaeger (2001a) dataset, where the encoding capacity of the model and readout modules are estimated through sequential recall of a delayed time-series signal under various time lags, quantified by the Pearson correlation between the target sequence and the model’s predicted memorization signal.
@@ 4.2 Experiments
The general procedure and pipeline for all network models are described in Section 3. Each model is evaluated on 7 experiments each according to the 7 readout Yeo networks, with the synthetic datasets identically generated for all functional networks on each of the same 1000 tests. The models are primarily evaluated on F1-score (for PDM, PDMDR, CDM) and Pearson correlation (for MemCap). Details on the experimental setup and implementation libraries are included in the Appendix B.
@@ 4.3 Empirical Performance: Processing and Memorization
We evaluate the performance of structural and null models under the mentioned processing and memorization synthetic tasks, showing the performance of each model over multiple reservoir spectral radius (alpha) in Figure 3, 4. Models under perceptual and memorization tasks show performance decay from $\alpha=1$, while contextual models perform well even in the chaotic $\alpha>1$ regime. The memory capacity of reservoirs shows the most notable decline in the chaotic regime, similar to results obtained in Suarez et al. (2020), and all models perform better than the fully connected bipartite null counterpart with the same read-in and read-out node count, showing gains in having higher topological complexity when strictly compared to 2-layer feedforward model. Model performance across 1000 different instantiations of the reservoir is fairly stable, showing that degree-conserving null models and bipartite model performed worse than the original biological structural model, structure-based subgraph model, and uniformly randomized structural weight model at close to criticality $\alpha=0.95$ (Figure 5, 6).
@@ 4.5 Comparison with other RNNs and testing with real-world data
The models tested follow the same notation described in prior work on COVID-19 prediction Panagopoulos et al. (2021). The additional models we tested are briefly described in Appendix B.4.
@@ 5 Functional sub-circuits’ topological performance Analysis
In this section, we analyze the neuro-physiological characteristics of a priori functional networks via two viewpoints: i) statics (e.g., size, betweenness, modularity) and dynamics (e.g., communicability) properties, quantifying their relationship with model performance Freeman (1977); Estrada & Hatano (2008); Clauset et al. (2004b).
@@ 5.1 Read-out functional networks’ statics properties
In this sub-section, we investigate FN statics properties through their node count statistics and betweenness measures Puzis et al. (2007). Based on Figure 5, we see that there exists an association, for some FNs, between their sizes (through node count statistics) and their betweenness scores. Specifically, DMN, the largest sub-circuits, also has the highest betweenness score. More importantly, there exists an association between the number of best performance analyses based on FN (Figure 4) to their betweenness scores. Specifically, DMN, ranked top 2 in 3 out of 4 tasks per Figure 4, ranked first in betweenness score. Other ”big” sub-circuits such as VIS or SM also have competitive betweenness.

### v2

@@ 1 Introduction
One of the prominent directions in building high-performance or general intelligence artificial systems is implementing biologically plausible models, particularly using human brain connectomes as the underlying graph of neural networks [1, 2, 3, 4]. Most state-of-the-art biologically inspired artificial connectomes are realized as reservoir networks with structural human brain data. However, structural connectomes (SC) derived from diffusion Magnetic Resonance Imaging (MRI) are generally rigid in both topology and scale compared to functional connectomes (FC) from functional MRI data. Thus, in this work, we explore the prospect of applying functional connectomes to reservoir Echo-state Networks (ESNs) under various conditions to comprehensively evaluate model performance, as well as build a consistent framework for implementations of connectomic data in artificial networks. Motivated by the fact that the topology of a neural network has a tangible impact on ESN performance, and that complex network topology sometimes improves the performance of the model, we systematically explore more complex and biologically inspired topology in reservoir neural networks. A critical step in ESN embedding requires pre-determined read-in and read-out layers with inputs sampled from the induced subgraphs of the embedded reservoir. Typically, a priori set of functional networks (FNs), such as Yeo’s FNs [5], is used as pre-determined subgraphs for read-in and read-out input selections. Since a priori set of FNs are derived from functional MRIs, yet, these pre-determined sub-circuits are mapped onto SC, it is unknown whether FN-to-SC embedding is well justified from i) neuro-physiological perspective and ii) ESN performance. In this paper, our contributions are as follows: i) We proposed a pipeline to implement and evaluate echo-state reservoir networks with various embedded topology on various processing and memorization tasks. To this end, we tested various topological embedding and evaluated their respective performance on diverse tasks, showing different performance optimums are highly dependent on the embedded neuro-physiological architecture of a priori functional networks and the corresponding task configurations; ii) we show that, in general, complex topology perform better than simple bipartite models or null models with feed-forward properties commonly seen in MLP for certain tasks and reservoir criticality conditions. As such, we also found that, contrary to earlier literature, the performance of the reservoir model is not strictly defined by the reservoir’s echo-state property defined by the spectral radius of the connectivity matrix; iii) we found that there exists a differentiated performance of functional networks induced from the human structural connectome with respect to different cognitive tasks and their corresponding computation-memorization balances. We analyzed the topological properties of a priori functional networks and highlighted their graph-theoretical properties that play significant roles in determining the final performance of the echo-state model.
@@ 2.1 Topologically-Embedded artificial neural network and Echo-State-Networks (ESN)
Echo-state networks are a variant of the classical reservoir computing (RC) paradigms, where the reservoir is a recurrent layer of the classical recurrent neural network [6]. RCs are typically proposed in place of regular RNN variants as a lightweight model, where the hidden states are high dimensions while only the readout layer weights of the model must be updated during training. The frozen nature of the reservoir layer allows any topology and unit dynamics to be embedded. As a result, RCs have been implemented as predictors of chaotic dynamics in physics simulators, or used as a time-series prediction model in general [7, 8, 9, 10, 11]. ESNs model time as discrete steps, making them suitable for discrete-time sampling data seen in real-world time-series datasets [12]. In the current literature, the relationship between the topological structure of an NN and its performance is still unclear. While complex networks may yield higher predictive performance or parameter efficiency [13], it is difficult to generalize randomly wired models outside the investigated context. Under the constraint of MLP-random interpolation, several random network models rewired similarly to the small-world regimes perform well on several real-life problems [14, 15]. Different topology perform well in limited tasks on small network sizes enabled by the learning matrix sequential algorithm for training directed acyclic graphs (DAGs). Random graph topology are also shown to perform well as image classifiers on the ImageNet dataset, especially the Watts-Strogatz (WS) model of small-world networks [16]. [17] echoes the statement on the performance of small-world models in the case of synthetic data, albeit the difference between WS and other complex graph families is less pronounced, possibly owing to the sufficient degree of topological complexity for function representation.
@@ 2.2 Brain Connectomics and human brain functional sub-circuits
Human Brain connectomics studies how the human brain is structurally connected [18, 19, 20, 21] and how a heterogeneous repertoire of resting-state or task-related functional circuits emerge on top of it [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 21, 32, 33]. There are two types of mesoscopic structures in large-scale brain networks: localized and non-localized. In this paper, we focus on investigating localized mesoscopic structures which are sub-systems learned from local/quasi-local network properties such as brain regions of interest (ROIs) or functional edges, or correlations among neighboring nodes [27]. In brain connectomics, these sub-structures are induced from a wide array of techniques, including but not limited to clustering [5, 34] or low-dimensional approximation of high-dimensional dynamics [35, 36, 37]. The most commonly known localized mesoscopic structures in brain networks are often referred to as functional sub-circuits or functional networks (FNs) [5, 38, 39, 27]. From here on, we will use the two terms functional sub-circuits and FNs interchangeably.
@@ 2.3 Human-Brain-Connectomics ESNs
In the context of modeling the human brain connectome, ESNs are implemented as prototype models of the structural brain topology. [2] shows that human brain connectivity captured through diffusion imaging performs better than random null network models in the critical regime, while also demonstrating computationally relevant properties of resting-state functional brain network parcellation. Given a minimum level of randomness and connection diversity from the original structural connectome, biologically inspired reservoir networks’ performance matches that of any classical random network model [1]. The authors of [40] conclude that the modular structure of reservoirs significantly impacts prediction error among multiple other features. While the connections are modeled in a recurrent network rather than strictly an RC, [41] shows that training an RNN constrained by reservoir features such as spatial wiring cost and neuron communicability leads to the emergence of structural and functional features commonly found in primate cerebral cortices.
@@ 3.1 Preliminary: Echo-state Implementation
To examine different topologies as computational graphs, we utilized the implementation of the classical ESN [42] from the conn2res package [4] which has the following form:
@@
where $\textbf{X}\in\mathds{R}^{N_{p}\times T}$ is the matrix containing temporal reservoir states $x[t]$ of a certain $N_{p}$ nodes ($N_{p}<N_{r}$) of the reservoir computed from the previous states and the input $u[t]$ for $t=1,...,T$, $\textbf{Y}\in\mathds{R}^{N_{o}\times T}$ is the ground-truth matrix containing $y[t]$, $\textbf{I}\in\mathds{R}^{N_{p}\times N_{p}}$ is the identity matrix, and $\lambda$ is the regularization parameter. Additionally, we keep track of the spectral radius $\alpha$ of the reservoir’s adjacency matrix to examine the reservoir’s adherence to the echo-state property (ESP), a condition quantifying the reservoir’s unique stable input-driven dynamics [42].
@@ 3.2 Pipeline and structural control model
The reservoir network is generally partitioned using intrinsic functional networks, a connectivity-based partition of brain network into functionally similar groups of areas [43]. The reservoir input weight matrix $\textbf{W}^{I}\in\mathds{R}^{n\times m}$ route the input signals to only certain $m$ nodes in the reservoir (Figure 1, 2). The input nodes are chosen as an entire functional network to examine message propagation on the reservoir computational graph; in this paper, we chose subcortical regions as input nodes, owning to the plausibility of these regions serving as relay stations for incoming sensory signals. Each of the other seven cortical regions is chosen as the read-out node-set, where the readout weight matrix $\textbf{W}^{O}$ connects the $N_{p}$ nodes of the functional network to the readout module.
Figure 2: Human brain functional networks (sub-circuits) as parcellated by Yeo and colleagues [5]. Figure is referenced from [44].
@@
We embed the structural human brain connectome into the reservoir as the default topology map. Consensus adjacency matrices are constructed from the original connectomes with bootstrap resampling to provide a reliable estimate of the model performance. The structural connectome is grouped into seven intrinsic functional networks and one subcortical region according to the Yeo functional mapping and parcellation [45]. These structures are known as sub-circuits or FNs, previously defined in Section 2.2.
@@ 3.3 Variants: Complex graph null models
We preserve the original degree-sequence of the structural connectome and rewired each edge in the connectome exactly once using the Maslov-Sneppen algorithm [46]. From the rewiring procedure, we obtained a reference null connectome with the original constraint still moderately enforced.
@@
An entirely new graph is constructed from the degree sequence of the original structural connectome, as opposed to only rewiring each edge once. The configuration model generates a random undirected pseudograph by randomly assigning edges to match a given degree sequence [47]. Self-loops are then removed from the existing graph to obtain the symmetric null connectome adjacency matrix for the reservoir.
@@ 4.1 Data
Adapted from structural consensus connectomes from [2]. The connectome is divided into 463 or 1015 approximately equally sized nodes. A group-consensus approach is adopted to mitigate reconstruction inconsistencies [48] and network measure sensitivities between different maps. Functional connectomes are obtained from a publicly available dataset (Derived Products from HCP-YA fMRI, [49]). Structural networks are taken from [2] where 66 participants were scanned using diffusion MRI under no task demands, divided into 463 or 1015 approximately equally sized nodes. The dataset contains fMRI time-series parcellated using the Schaefer parcellation at resolutions from 100 to 900 parcels; we use the parcellation resolution at 500 nodes via the Yeo 7-region functional networks schema without Global Signal Regression (GSR) under resting state.
@@
We used synthetic behavioral data to simulate computation from multiple signal sources and assess the memory capacity of the reservoir under various means. The PerceptualDecisionMaking (PDM) [50] task is a two-alternative forced choice task requiring the model to integrate and compare the average of two stimuli, with the PerceptualDecisionMakingDelayResponse (PDMDR) [51] variant artificially inducing a memory requirement by prompting the model to answer after a random delay. The context-dependent decision-making task ContextDecisionMaking (CDM) [52] requires the model to make a perceptual decision based on only one of the two stimulus inputs from two different modalities, where a rule signal indicates the relevant modality. The assessment for the memory capacity of the model is the MemoryCapacity (MemCap) [53] dataset, where the encoding capacity of the model and readout modules are estimated through sequential recall of a delayed time-series signal under various time lags, quantified by the Pearson correlation between the target sequence and the model’s predicted memorization signal.
@@ 4.2 Experiments
The general procedure and pipeline for all network models are described in Section 3. Each model is evaluated on 7 experiments each according to the 7 readout Yeo networks, with the synthetic datasets identically generated for all functional networks on each of the same 1000 tests. The models are primarily evaluated on F1-score (for PDM, PDMDR, CDM) and Pearson correlation (for MemCap). Details on the experimental setup and implementation libraries are included in the Appendix LABEL:sec:exp_settings.
@@ 4.3 Empirical Performance: Processing and Memorization
We evaluate the performance of structural and null models under the mentioned processing and memorization synthetic tasks, showing the performance of each model over multiple reservoir spectral radius (alpha) in Figure 3, 4. Models under perceptual and memorization tasks show performance decay from $\alpha=1$, while contextual models perform well even in the chaotic $\alpha>1$ regime. The memory capacity of reservoirs shows the most notable decline in the chaotic regime, similar to results obtained in [2], and all models perform better than the fully connected bipartite null counterpart with the same read-in and read-out node count, showing gains in having higher topological complexity when strictly compared to 2-layer feedforward model. Model performance across 1000 different instantiations of the reservoir is fairly stable, showing that degree-conserving null models and bipartite model performed worse than the original biological structural model, structure-based subgraph model, and uniformly randomized structural weight model at close to criticality $\alpha=0.95$ (Figure 5, 6).
@@ 4.5 Comparison with other RNNs and testing with real-world data
The models tested follow the same notation described in prior work on COVID-19 prediction [54]. The additional models we tested are briefly described in Appendix LABEL:sec:covid_models.
@@ 5 Functional sub-circuits’ topological performance Analysis
In this section, we analyze the neuro-physiological characteristics of a priori functional networks via two viewpoints: i) statics (e.g., size, betweenness, modularity) and dynamics (e.g., communicability) properties, quantifying their relationship with model performance [55, 56, 57].
@@ 5.1 Read-out functional networks’ statics properties
In this sub-section, we investigate FN statics properties through their node count statistics and betweenness measures [58]. Based on Figure 5, we see that there exists an association, for some FNs, between their sizes (through node count statistics) and their betweenness scores. Specifically, DMN, the largest sub-circuits, also has the highest betweenness score. More importantly, there exists an association between the number of best performance analyses based on FN (Figure 4) to their betweenness scores. Specifically, DMN, ranked top 2 in 3 out of 4 tasks per Figure 4, ranked first in betweenness score. Other "big" sub-circuits such as VIS or SM also have competitive betweenness.